OpenStudy (anonymous):

3x-2x =-16 rewrite it in spole intercept form

OpenStudy (saifoo.khan):

is your equation correct? there is no "y".

OpenStudy (anonymous):

its in standard form and i need to change it to slope intercept form

OpenStudy (saifoo.khan):

3x-2x=-16 ??? check it

OpenStudy (anonymous):

oh my bad it is 3x-2y=-16

OpenStudy (saifoo.khan):

\[y = \frac32x+8\]

OpenStudy (anonymous):

cool but how did you find the answer?

OpenStudy (saifoo.khan):

3x-2y=-16 -2y = -3x - 16 2y = 3x + 16 \[y = \frac{3x+16}{2}\] \[y = \frac32x + \frac{16}{2}\] \[y = \frac32x + 8\]

OpenStudy (anonymous):

thanks

OpenStudy (saifoo.khan):

uw